OpenStudy (anonymous):

what is the cost to remove 75% of the pollutant? 6,000 C= --------- 200-2x using this formula how do you solve? I have a hard time with percentages

OpenStudy (anonymous):

sub in 3/4 for x it looks like so as long x is the variable for the polutant amount. 3/4 = 75%. If they give you another amount for the polutant say a cubic volume = 16, then multiply 3/4 times that and then put the answer in for x to get the cost. Otherwise 1 = 100%, and 3/4 = 75% for x. Let me know if you've got something else going on with it or you've got a question about it.

OpenStudy (anonymous):

sub in 3/4 for x it looks like so as long x is the variable for the polutant amount. 3/4 = 75%. If they give you another amount for the polutant say a cubic volume = 16, then multiply 3/4 times that and then put the answer in for x to get the cost. Otherwise 1 = 100%, and 3/4 = 75% for x. Let me know if you've got something else going on with it or you've got a question about it.